show that if the normals at the points P(at1 ,2at1 ) and G(at22 ,2at2 ) to the parabola y2 =4ax, intersect on the paarabola, then t1t2 = 2.

Grade:12th pass

1 Answers

Arun
6 years ago
 

Let the normals at P and Q meet at R(at2, 2at).

Then t = – t1 – 2/t1 and t = – t2 – 2/t2 .

Therefore t1 + 2/t1 = t2 + 2/t2 ⇒ (t1 – t2) = 2(t1-t2 )/(t1 t2 ) ⇒ t1t2 = 2.